# Attribute or Abstain: Large Language Models as Long Document Assistants

### v1

@@ Abstract
LLMs can help humans working with long documents, but are known to hallucinate. Attribution can increase trust in LLM responses: The LLM provides evidence that supports its response, which enhances verifiability. Existing approaches to attribution have only been evaluated in RAG settings, where the initial retrieval confounds LLM performance. This is crucially different from the long document setting, where retrieval is not needed, but could help. Thus, a long document specific evaluation of attribution is missing. To fill this gap, we present LAB, a benchmark of 6 diverse long document tasks with attribution, and experiment with different approaches to attribution on 4 LLMs of different sizes, both prompted and fine-tuned.
We find that citation, i.e. response generation and evidence extraction in one step, mostly performs best. We investigate whether the “Lost in the Middle” phenomenon exists for attribution, but do not find this. We also find that evidence quality can predict response quality on datasets with simple responses, but not so for complex responses, as models struggle with providing evidence for complex claims. We release code and data for further investigation [^1] Github repository, code under Apache 2.0, dataset licenses depend on original license, see §B..
@@ 1 Introduction
Recent LLMs can process long documents Shaham et al. (2023); Li et al. (2023b), showing great potential as long document assistants. For example (Fig. 1), such an assistant could answer a researcher’s questions about a paper. However, due to LLM hallucinations Slobodkin et al. (2023), the researcher must verify responses, which is difficult with lengthy papers. To improve verifiability and trust, the assistant should either attribute Rashkin et al. (2023) or abstain Slobodkin et al. (2023): If it finds the necessary information, it should provide a response and point to the evidence in the paper (attribute). If not, it should clearly communicate this (abstain). We investigate the capabilities of LLMs to fulfill these requirements, and the relation between response quality and evidence quality.
Formally, we assume an instruction $I$ (e.g. a question) and a document $D$ consisting of segments $d$ (e.g. paragraphs) [^2] In other scenarios, $D$ can be a large corpus, see below.. There are two subtasks that can be solved jointly or independently: One is to generate a response $R$ (e.g. an answer) containing statements $r$. If the response is not abstained (e.g. by saying “unanswerable”), the other subtask is to retrieve evidence $E_{i}\subset D$ for each $r_{i}$ such that $r_{i}$, is attributable to $E_{i}$, i.e. "according to $E_{i}$, $r_{i}$" is true Rashkin et al. (2023).
To decrease the input length in post-hoc and citation, $D$ can be reduced to $D^{\prime}\subset D$ via and additional initial retrieval step. This results in two further attribution approaches: (4) reduced-post-hoc and (5) reduced-citation.
Attribution has only been investigated in retrieval augmented generation (RAG) settings, where $D$ is not a document, but a large corpus (e.g. Wikipedia) that does not fit the LLM context. This means that only approaches with initial retrieval, retrieve-then-read, reduced-post-hoc and reduced-citation, can be used. Performance depends on retrieval quality, and the best-performing approach is unclear Bohnet et al. (2022); Gao et al. (2023b); Malaviya et al. (2024).
In contrast, when $D$ is a long document that fits the LLM context, the confounding initial retrieval can be omitted. Still, it is possible that separating generation and retrieval improves performance, as shown in related works on task decomposition Sun et al. (2023) and reducing long LLM inputs, where Agrawal et al. (2024) found positive effects, while Xu et al. (2024) and Bai et al. (2023) did not. These works did not consider attribution, so there is a lack of knowledge on the effect of separating response generation and evidence retrieval, and the optimal approach to attribution on long documents.
@@
Evidence quality (attributability) can be evaluated without external reference Honovich et al. (2022); Yue et al. (2023); Tang et al. (2024). If evidence quality were positively correlated with response quality, bad responses could be abstained from by filtering responses with bad evidence quality. If not, this would lead to abstaining from potentially helpful responses with insufficient evidence. Current research on the relation of response quality and evidence quality is inconclusive: Bohnet et al. (2022) and Gao et al. (2023b) reranked multiple sampled responses by evidence quality. While the former found an improvement in response quality, the latter did not. Neither provide an analysis, so we lack understanding if and how evidence quality correlates with response quality.
To close these gaps, we compile LAB, a Long-document Attribution Benchmark of 6 long-document datasets with diverse tasks (QA, classification, NLI and summarization) and domains (science, law, governmental and Wikipedia). We conduct experiments using the outlined attribution approaches on 4 LLMs of varying sizes, prompted and fine-tuned, to answer three research questions:
RQ1: What are optimal approaches for attribution in long document tasks? We find that large and fine-tuned models reach best evidence quality via citation, but small prompted LLMs can benefit from post-hoc evidence retrieval.
@@ 2 Related Work
Current research in attribution is done in three strains: First, some works evaluate post-hoc, retrieve-then-read and citation approaches in their ability to produce attributed responses Bohnet et al. (2022); Liu et al. (2023), some proposing new datasets Malaviya et al. (2024) or benchmarks DeYoung et al. (2020a); Gao et al. (2023b). Second, methodological works propose new fine-tuning Schimanski et al. (2024); Huang et al. (2024) and prompting Berchansky et al. (2024); Fierro et al. (2024) methods to improve the citation capabilities of language models, but do not compare to retrieve-then-read or post-hoc approaches. Both of these strains have focused on open domain QA, and neglected the long document scenario. Here, we close these gaps by providing a comprehensive investigation of attribution for long documents.
To evaluate attributability automatically, most works have used TRUE, Flan-T5-XXL fine-tuned on several NLI datasets Honovich et al. (2022); Bohnet et al. (2022); Gao et al. (2023b); Fierro et al. (2024); Huang et al. (2024). More recently, Attrscore Yue et al. (2023) and Minicheck Tang et al. (2024) were proposed specifically for the evaluation of attributability. We compare these models and employ the best-performing for attributability evaluation.
@@ 3.1 Datasets
Table 1: The datasets in LAB span multiple domains and task types. The numbers for Natural Questions differ from the original publication, as we filtered the instances (§B). Column names: Doc #: Average number of words per document. Res $|R|$/#W: Number of statements / Average number of words per response. #Evi: Number of annotated evidence segments per instance. Evi Lvl: Level of annotated evidence. Tasks: QA: Question answering. Cls: Classification. FC: Fact checking. NLI: Natural Language Inference. Sum: Summarization. [1] Dasigi et al. (2021) [2] Kwiatkowski et al. (2019) [3] DeYoung et al. (2020b) [4] Kamoi et al. (2023) [5] Koreeda and Manning (2021) [6] Huang et al. (2021)
The datasets in LAB are shown in Table 1. All datasets are in English. GovReport Huang et al. (2021) is the only dataset without annotated evidence. To simulate gold evidence, we used BM25 to find the 2 best-matching paragraphs from a document for each sentence in the gold summary similar to Ravaut et al. (2024). Due to limited resources, we use at most 2000 test instances from any dataset (100 for GPT-4). For details and examples see §B.
@@ 3.2 Evaluation
Response parsing
For all datasets except GovReport, responses consist of single statements. For GovReport, we split responses into statements (sentences) using NLTK Bird (2006). For citation, LLMs are expected to generate segment identifiers (“[1] [2]”) at the end of each statement.
Evaluation of response quality
For comparability with related work, we used established metrics to evaluate response quality: Exact match F1 (EM) [^5] Found to correlate better than BERTScore Zhang* et al. (2020) with human judgment in our preliminary experiments. on QASPER Dasigi et al. (2021); Shaham et al. (2023) and Natural Questions Kwiatkowski et al. (2019); Bohnet et al. (2022), classification macro F1 (CF1) for Evidence Inference DeYoung et al. (2020b), Wice Kamoi et al. (2023) and ContractNLI Koreeda and Manning (2021) and ROUGE-L [^6] ROUGE has shown good correlation with human relevance judgments in Wu et al. (2024). (RL, Lin 2004) for GovReport Huang et al. (2021); Shaham et al. (2023).
@@
For the other datasets, evidence F1 is insufficient: GovReport does not come with annotated evidence, and for datasets with free-form responses (QASPER and Natural Questions), evidence F1 is too rigid: A model might produce a response different from the ground truth, but supported by the retrieved evidence. Even though evidence quality is high, evidence F1 might be low. For these datasets, we evaluate evidence quality as attributability [^7] Also known as citation recall Gao et al. (2023b)., Gao et al. (2023b); Huang et al. (2021); Schimanski et al. (2024). We assume an attributability evaluation model $\mathrm{M_{a}}(E,r)\rightarrow\{1,0\}$. The attributability score (AS) of a response is computed as the proportion of attributable statements.
@@
When a response is abstained, we do not evaluate attributability, as no evidence is required. We additionally evaluate Unanswerable F1 for QA datasets:
Unanswerable F1 (UF1)
To detect abstaining, we compiled a list of keywords based on Slobodkin et al. (2023) and manual inspection. If a response contained a keyword, any predicted evidence was removed, and the response was set to "unanswerable" (see §G).
@@ 3.3 Generation
We focus on two groups of LLMs with at least 8K tokens input length: (1) The large state of the art models GPT-3.5888gpt-35-turbo-0613-16k and GPT-4999gpt-4-turbo-128k, as they hold top positions in other long document benchmarks Li et al. (2023b); Shaham et al. (2023); An et al. (2023) (2) Small (~3-7B) models that are accessible with limited resources. We tested a range of models with citation on QASPER and GovReport and selected Longchat-7B and Flan-T5-XL as the best performing for prompting and fine-tuning, respectively. For complete selection results and hyperparameters see §E.
@@ 3.4 Retrieval
In retrieve-then-read and reduced, queries were constructed based on information available in the instruction (e.g. question or a claim). For GovReport, similar to Zhang et al. (2020), we created queries from all document segments and retrieved paragraphs based on self-similariy. In post-hoc, queries were constructed based on the instruction and the generated response. For both post-hoc and retrieve-then-read, we set $k=5$ for Wice, and $k=2$ for all other datasets based on evidence statistics (see Table 1). In reduced approaches we set $k=10$ based on Xu et al. (2024) (see §F).
@@ 4 Experiments
QSP NQ EI WIC CNLI GR Average EM ATT UF1 EM ATT UF1 CF1 EF1 CF1 EF1 CF1 EF1 RL ATT RQ EQ GPT-3.5 p-h 45 62 65 51 44 58 77 22 52 48 44 40 26 74 49.17 48.33 rtr 42 78 51 50 42 57 73 25 50 44 47 41 21 40 47.17 45.00 cit 52 60 67 47 38 53 78 50 52 59 43 53 26 59 49.67 53.17 r-p-h 48 71 58 50 45 57 78 22 52 48 45 40 22 64 49.17 48.33 r-cit 50 70 64 46 38 53 78 52 48 63 48 57 22 58 48.67 56.33 GPT-4 p-h 65 68 68 52 42 56 87 24 32 36 66 50 27 63 54.83 47.17 rtr 56 83 56 58 36 61 74 31 22 28 64 59 20 27 49.00 44.00 cit 68 76 71 51 49 57 86 64 35 47 63 64 27 62 55.00 60.33 r-p-h 62 71 69 51 43 57 84 24 28 30 65 50 22 54 52.00 45.33 r-cit 63 76 66 54 49 58 83 49 33 46 64 67 22 56 53.17 57.17 Flan-T5 p-h 55 67 58 81 61 74 86 22 44 46 77 57 27 90 61.67 57.17 rtr 43 74 55 79 62 74 77 25 43 43 64 58 19 67 54.17 54.83 cit 53 57 56 84 75 71 83 59 53 70 77 78 25 71 62.50 68.33 r-p-h 53 72 58 80 62 75 87 22 45 46 76 58 23 84 60.67 57.33 r-cit 52 62 55 84 75 73 84 57 53 68 75 75 22 68 61.67 67.50 Longchat p-h 23 57 62 21 33 38 66 22 23 31 38 33 24 67 32.50 40.50 rtr 29 56 49 27 32 43 43 25 19 25 39 39 21 41 29.67 36.33 cit 22 53 52 17 2 38 66 13 24 19 36 14 23 9 31.33 18.33 r-p-h 33 58 54 25 35 41 61 22 20 27 33 30 22 63 32.33 39.17 r-cit 26 48 53 21 4 42 65 13 20 25 24 12 22 24 29.67 21.00

### v2

@@ Abstract
LLMs can help humans working with long documents, but are known to hallucinate. Attribution can increase trust in LLM responses: The LLM provides evidence that supports its response, which enhances verifiability. Existing approaches to attribution have only been evaluated in RAG settings, where the initial retrieval confounds LLM performance. This is crucially different from the long document setting, where retrieval is not needed, but could help. Thus, a long document specific evaluation of attribution is missing. To fill this gap, we present LAB, a benchmark of 6 diverse long document tasks with attribution, and experiments with different approaches to attribution on 5 LLMs of different sizes.
We find that citation, i.e. response generation and evidence extraction in one step, performs best for large and fine-tuned models, while additional retrieval can help for small, prompted models. We investigate whether the “Lost in the Middle” phenomenon exists for attribution, but do not find this. We also find that evidence quality can predict response quality on datasets with simple responses, but not so for complex responses, as models struggle with providing evidence for complex claims. We release code and data for further investigation [^1] Github repository, code under Apache 2.0, dataset licenses depend on original license, see §B..
@@ 1 Introduction
Recent LLMs can process long documents Shaham et al. (2023); Li et al. (2023b), showing great potential as long document assistants. For example (Fig. 1), such an assistant could answer a researcher’s questions about a paper. However, due to LLM hallucinations Slobodkin et al. (2023), the researcher must verify responses, which is difficult with lengthy papers. To improve verifiability and trust, the assistant should either attribute Rashkin et al. (2023) or abstain Slobodkin et al. (2023): If it finds the necessary information, it should provide a response and point to the evidence in the paper (attribute). If not, it should clearly communicate this (abstain). We investigate the capabilities of LLMs to fulfill these requirements, and the relation between response quality (i.e. correctness) and evidence quality (i.e. the relevance of the evidence to the response).
Thus far, attribution has only been investigated in retrieval augmented generation (RAG) settings, where evidence comes from a large corpus (e.g. Wikipedia) that does not fit the LLM context. This means that some form of retrieval is required to provide the LLM with a limited number of relevant passages or to retrieve evidence post-hoc. Performance depends on retrieval quality, and the best-performing approach is unclear Bohnet et al. (2022); Gao et al. (2023b); Malaviya et al. (2023).
In contrast, when the potential evidence is a single long document that fits the LLM context window, the confounding retrieval can be omitted, as LLMs can cite from their input Gao et al. (2023b). Still, it is possible that separating response generation and evidence retrieval improves performance, as shown in related works on task decomposition Sun et al. (2023) and reducing long LLM inputs, where Agrawal et al. (2024) found positive effects, while Xu et al. (2024) and Bai et al. (2023) did not. These works did not consider attribution, so there is a lack of knowledge on the effect of separating response generation and evidence retrieval, and the optimal approach to attribution on long documents.
@@
Evidence quality can be automatically evaluated without external reference Honovich et al. (2022); Yue et al. (2023); Tang et al. (2024). If evidence quality were positively correlated with response quality, bad responses could be filtered by identifying responses with bad evidence quality, thereby improving abstaining. If not, this would lead to abstaining from potentially helpful responses with insufficient evidence. Current research on the relation of response quality and evidence quality is inconclusive: Bohnet et al. (2022) and Gao et al. (2023b) reranked multiple sampled responses by evidence quality. While the former found an improvement in response quality, the latter did not. Neither provide an analysis, so we lack understanding if and how evidence quality correlates with response quality.
To close these gaps, we compile LAB, a Long-document Attribution Benchmark of 6 long-document datasets with diverse tasks (QA, classification, fact checking, NLI and summarization) and domains (science, law, governmental and Wikipedia). We conduct experiments using approaches to attribution with and without additional retrieval on 5 LLMs of varying sizes, prompted and fine-tuned, to answer three research questions:
RQ1: What are optimal approaches for attribution in long document tasks? We find that large and fine-tuned models reach best evidence quality by directly citing from input documents, but small prompted LLMs can benefit from post-hoc evidence retrieval.
@@ 2 Related Work
Current research in attribution is done in three strains: First, some works evaluate a range of approaches in their ability to produce attributed responses Bohnet et al. (2022); Liu et al. (2023), some proposing new datasets Malaviya et al. (2023) or benchmarks DeYoung et al. (2020a); Gao et al. (2023b). Second, methodological works propose new fine-tuning Schimanski et al. (2024); Huang et al. (2024) and prompting Berchansky et al. (2024); Fierro et al. (2024) methods to improve the citation capabilities of language models, but do not compare to approaches with additional retrieval. Both of these strains have focused on open domain QA, and neglected the long document scenario. Here, we close these gaps by providing a comprehensive investigation of attribution for long documents.
To evaluate evidence quality automatically, most works have used TRUE, Flan-T5-XXL fine-tuned on several NLI datasets Honovich et al. (2022); Bohnet et al. (2022); Gao et al. (2023b); Fierro et al. (2024); Huang et al. (2024). More recently, Attrscore Yue et al. (2023) and Minicheck Tang et al. (2024) were proposed specifically for the evaluation of attributability. We compare these models and employ the best-performing for attributability evaluation.
@@
3 Formalization
We first define the task of producing attributed responses. Based on this, we define the approaches to attribution compared in this work.
Task definition
We assume an instruction $I$ (e.g. a question) and a document $D$ consisting of segments $d$ (e.g. paragraphs). There are two subtasks that can be solved jointly or independently: One is to generate a response $R$ (e.g. an answer) containing statements $r$. If the response is not abstained (e.g. by saying “unanswerable”), the other subtask is to retrieve evidence $E_{i}\subset D$ for each $r_{i}$ such that $r_{i}$, is attributable to $E_{i}$, i.e. "according to $E_{i}$, $r_{i}$" is true Rashkin et al. (2023).
Approaches to attribution
To decrease the input length in post-hoc and citation, $D$ can be reduced to $D^{\prime}\subset D$ via an additional initial retrieval step. This results in two further attribution approaches: (4) reduced-post-hoc and (5) reduced-citation.
@@ 4 Methods
We first introduce the datasets (§4.1) and evaluation metrics (§4.2) used in LAB, our long document attribution benchmark. We then describe our experimental setup for generation (§4.3) and retrieval (§4.4).
@@ 4.1 Datasets
Table 1: The datasets in LAB span multiple domains and task types. The numbers for Natural Questions differ from the original publication, as we filtered the instances (§B). Column names: Doc #W: Average number of words per document. Res $|R|$/#W: Number of statements / Average number of words per response. #Evi: Number of annotated evidence segments per instance. Evi Lvl: Level of annotated evidence. Tasks: QA: Question answering. Cls: Classification. FC: Fact checking. NLI: Natural Language Inference. Sum: Summarization. [1] Dasigi et al. (2021) [2] Kwiatkowski et al. (2019) [3] DeYoung et al. (2020b) [4] Kamoi et al. (2023) [5] Koreeda and Manning (2021) [6] Huang et al. (2021)
The datasets in LAB are shown in Table 1. All datasets are in English. GovReport Huang et al. (2021) is the only dataset without annotated gold evidence. To simulate gold evidence, we use BM25 to find the 2 best-matching paragraphs from a document for each sentence in the gold summary similar to Ravaut et al. (2023). Due to limited resources, we use at most 2000 test instances from any dataset (100 for GPT-4). For details and examples see §B.
@@ 4.2 Evaluation
We evaluate two main properties of attributed responses: 1. Response quality, i.e. whether a model-generated response is correct in reference to an annotated ground-truth. 2. Evidence quality, i.e. whether the provided evidence is relevant to the given response. These are evaluated with task-specific metrics (see below). For QA datasets, we additionally evaluate Unanswerable F1, i.e. whether models abstain from responding to unanswerable instances.
4.2.1 Response Quality
For comparability with related work, we used established metrics to evaluate response quality: Exact match F1 (EM) [^4] Found to correlate better than BERTScore Zhang et al. (2020b) with human judgment in our preliminary experiments. on QASPER Dasigi et al. (2021); Shaham et al. (2023) and Natural Questions Kwiatkowski et al. (2019); Bohnet et al. (2022), classification macro F1 (CF1) for Evidence Inference DeYoung et al. (2020b), Wice Kamoi et al. (2023) and ContractNLI Koreeda and Manning (2021) and ROUGE-L [^5] ROUGE has shown good correlation with human relevance judgments in Wu et al. (2024). (RL, Lin 2004) for GovReport Huang et al. (2021); Shaham et al. (2023).
4.2.2 Evidence Quality
@@
For the other datasets, evidence F1 is insufficient: GovReport does not come with annotated evidence, and for datasets with free-form responses (QASPER and Natural Questions), evidence F1 is too rigid: A model might produce a response different from the ground truth, but supported by the retrieved evidence. Even though the predicted evidence is relevant to the predicted response, evidence F1 might be low. For these datasets, we evaluate evidence quality as attributability [^6] Also known as citation recall Gao et al. (2023b). Rashkin et al. (2023), Gao et al. (2023b); Huang et al. (2021); Schimanski et al. (2024). We assume an attributability evaluation model $\mathrm{M_{a}}(E,r)\rightarrow\{1,0\}$. The attributability score (AS) of a response is computed as the proportion of attributable statements.
@@
On QA tasks, models can abstain from responding. In these cases, we do not evaluate attributability, as no evidence is required. Because this could encourage abstaining from responding too often, we additionally evaluate Unanswerable F1 (see below).
@@ Results
4.2.4 Unanswerable F1 (UF1)
To detect abstaining, we compiled a list of keyphrases based on Slobodkin et al. (2023) and manual inspection. If a response contained a keyphrase, any predicted evidence was removed, and the response was set to "unanswerable" (see §G).
@@ 4.3 Generation
We focus on two groups of LLMs with at least 8K tokens input length: (1) The large state of the art models GPT-3.5777gpt-35-turbo-0613-16k and GPT-4888gpt-4-turbo-128k, as they hold top positions in other long document benchmarks Li et al. (2023b); Shaham et al. (2023); An et al. (2024) (2) Small (~3-7B) models that are accessible with limited resources. We tested a range of models with citation on QASPER and GovReport and selected Longchat [^9] Longchat-7B-v1.5-32K Li et al. (2023a) and Mistral [^10] Mistral-7B-Instruct-v0.2 Jiang et al. (2023) as the best-performing for prompting and Flan-T5  [^11] Flan-T5-XL Longpre et al. (2023) as the best performing for fine-tuning. For complete selection results and hyperparameters see §E.
@@
Response parsing
For all datasets except GovReport, responses consist of single statements. For GovReport, we split responses into statements (sentences) using NLTK Loper and Bird (2002). For citation, LLMs are expected to generate segment identifiers (“[1] [2]”) at the end of each statement.
@@ 4.4 Retrieval
In retrieve-then-read and reduced, queries were constructed based on information available in the instruction (e.g. question or a claim). For GovReport, similar to Zhang et al. (2020a), we created queries from all document segments and retrieved paragraphs based on self-similarity. In post-hoc, queries were constructed based on the instruction and the generated response. For both post-hoc and retrieve-then-read, we retrieve 5 evidence segments for Wice, and 2 for all other datasets based on evidence statistics (see Table 1). In reduced approaches we reduce input documents to 10 segments based on Xu et al. (2024) (see §F).
@@ 5 Experiments
QSP NQ EI WIC CNLI GR Avg EM ATT UF1 EM ATT UF1 CF1 EF1 CF1 EF1 CF1 EF1 RL ATT RQ EQ GPT-3.5 p-h 45 62 65 51 44 58 77 22 52 48 44 40 26 74 49.17 48.33 rtr 42 78 51 50 42 57 73 25 50 44 47 41 21 40 47.17 45.00 cit 51 55 70 47 38 53 78 50 52 59 43 53 26 59 49.50 52.33 r-p-h 48 71 58 50 45 57 78 22 52 48 45 40 22 64 49.17 48.33 r-cit 50 69 64 46 38 53 78 52 48 63 48 57 22 58 48.67 56.17 GPT-4 p-h 65 68 68 52 42 56 87 24 32 36 66 50 27 63 54.83 47.17 rtr 56 83 56 58 36 61 74 31 22 28 64 59 20 27 49.00 44.00 cit 68 76 71 51 49 57 86 64 35 47 63 64 27 62 55.00 60.33 r-p-h 62 71 69 51 43 57 84 24 28 30 65 50 22 54 52.00 45.33 r-cit 63 76 66 54 49 58 83 49 33 46 64 67 22 55 53.17 57.00 Flan-T5 p-h 55 67 58 81 61 74 86 22 44 46 77 57 27 90 61.67 57.17 rtr 43 74 55 79 62 74 77 25 43 43 64 58 19 67 54.17 54.83 cit 53 57 56 84 75 71 83 59 53 70 77 78 25 71 62.50 68.33 r-p-h 53 72 58 80 62 75 87 22 45 46 76 58 23 84 60.67 57.33 r-cit 52 62 55 84 75 73 84 57 53 68 75 75 22 68 61.67 67.50 Longchat p-h 23 57 62 21 33 38 66 22 23 31 38 33 24 67 32.50 40.50 rtr 29 56 49 27 32 43 43 25 19 25 39 39 21 41 29.67 36.33 cit 21 6 60 17 2 38 66 13 24 19 36 14 25 9 31.50 10.50 r-p-h 33 58 54 25 35 41 61 22 20 27 33 30 22 63 32.33 39.17 r-cit 26 26 53 21 4 42 65 13 20 25 24 12 22 23 29.67 17.17 Mistral p-h 32 67 59 28 42 43 75 20 37 45 41 40 22 63 39.17 46.17 rtr 27 76 50 30 42 46 66 25 37 42 51 43 20 39 38.50 44.50 cit 35 39 61 26 15 41 75 27 34 39 47 38 22 0 39.83 26.33 r-p-h 31 71 58 29 44 43 76 21 38 45 45 38 21 64 40.00 47.17 r-cit 36 53 58 28 21 42 77 26 35 47 51 44 22 1 41.50 32.00